# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ 4.3 Outflow Formation and Blowout in a Starburst Galaxy
It may be tempting to conclude that hot gas or cosmic rays from these supernovae are driving the shells’ expansion, especially since, for the lowest velocity shell at 169 $\rm km~s^{-1}$, the dynamical time to traverse the inner 500 pc is $\la 4$ Myr, well after the onset of CCSNe for a burst at least 10 Myr old. That being said, the age reported by Parker2025 from fits to the UV continuum that Mrk 1486 corresponds to a combination of a substantial young ($\leq 5$ Myr) stellar population and a far older (40 Myr) population. In light of the more detailed star formation history, while CCSNe may be the progenitors of the material in these expanding shells, they may not necessarily be responsible for their acceleration. Here, the shape of the velocity field from OutLines can provide insight into the driving mechanism. Indeed, the velocity index of $\beta=0.7\pm 0.2$ disfavors (but cannot fully not reject) the rapid acceleration associated with ram pressure driving ($\beta\la 0.5$, e.g., 2005ApJ...618..569M; 2016MNRAS.463..541C). However, these $\alpha$-rich shells are more likely continuously accelerated by optically thin line-driven radiation pressure associated with $\beta=0.8$ (e.g., 1986A&A...164...86P). Moreover, the total energy injection rate of $10^{40}\rm~erg~s^{-1}$ for the series of shells is roughly consistent with stellar winds, which have $L_{mech}\approx 10^{39}\rm~erg~s^{-1}$ for a given stellar population (see, e.g., Starburst99 or BPASS, starburst99; Byrne2022, respectively). Radiative driving by young stellar populations may also account for the less pronounced C ii troughs via ionizing feedback. Furthermore, given that the travel time across the central kpc is $\la 4$ Myr, the >40>40 Myr stellar populations are unlikely to be responsible for these superbubbles. Corroborating evidence for optically thick radiation pressure is the presence of the fluorescent $\lambda 1194,1197$ lines but the absence of resonant line infilling. Such a scenario occurs in the optically thick line limit where photon escape primarily occurs through the fluorescent channel (e.g., 2015ApJ...801...43S, their Figures 5-6).
@@ 4.4 Thermal Winds in AGN Radio Lobes
The Seyfert 2 galaxy NGC 2992 has long been known to exhibit conical outflows (Allen1999; Veilleux2001). Using archival observations from the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) made using WiFES (WiFES), we have confirmed broad, asymmetric [O iii] $\lambda\lambda 4959,5007$ lines after summing over all the spaxels to boost signal to noise in the line profile and to mitigate any aperture effects. The low resolution ($R\approx$ 3 000 or FWHM of 105 $\rm km~s^{-1}$) requires corrections for the LSF by convolving with a Gaussian kernel with $\sigma_{\rm LSF}=$45 $\rm km~s^{-1}$; however, we note that the LSF predominantly impacts the narrow Gaussian core. The high signal in the wings ($S/N\approx 3$ in the blue broad extent of [O iii] $\lambda 4959$) might allow for tight constraints on the wind properties even without LSF corrections – indeed fitting with and without LSF corrections yields similar results for the OutLines parameters. Nevertheless, incorporating the LSF into the forward modeling with OutLines is essential to remain consistent with the observations. As a part of the LSF correction, we fold in the summed best-fit stellar continuum included in S7 to account for stellar contributions to the continuum, including any contamination of the line profile.
@@
The bright intermediate portion of the wings suggests the acceleration power law for the velocity field. Visual assessment of the highly asymmetric line profile in the spectrum of NGC 2992 and comparison with models in Figures 11, 12, and 13 also suggests a hollow cone geometry with an obstructing disk. As such, we subsequently model the [O iii] doublet using OutLines assuming an acceleration power law, an open cone geometry, and an obstructing disk. We impose the following into the prior: $v_{\infty}$ within a 300 $\rm km~s^{-1}$ standard deviation of the apparent maximum velocity of 1100 $\rm km~s^{-1}$, a static velocity dispersion within 50 $\rm km~s^{-1}$ standard deviation of the apparent dispersion of 110 $\rm km~s^{-1}$, accounting for the LSF, line fluxes within 5% of the [O iii] 5007/4959 flux ratio of 2.98 given by atomic data (Storey2000), and angles within 10 standard deviation of the best guess values from the trust-reflective fit. With line profiles predicted by OutLines added to the continuum and subsequently convolved with the WiFES LSF (WiFES), we use emcee to sample the posterior on the outflow parameters, requiring $\theta_{o}-\theta_{c}>5^{\circ}$. Due to the large number of parameters, potential degeneracies, and sensitivity to initial conditions, we run the sampler four times longer, for a total of 40 000 samples post burn-in. Implementation of our modeling is demonstrated for reference in Appendix D with the best-fit profile shown in Figure 19 and results listed in Table 1. We note that the posterior sampling ultimately rejects the priors on the angles as the best-guess parameters fell within a local minimum.
@@
Figure 19: MCMC fit of an OutLines.Nebular model to the [O iii] $\lambda\lambda 4959,5007$ doublet observed in the nucleus of Seyfert galaxy NGC 2992 as part of the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, S7DR2) DR2, assuming a hollow cone plus disk outflow geometry. Line styles and colors as in Figure 16. Blue line indicates the profile after LSF correction.
@@
The slope $\beta=2.48$ indicates, for the acceleration power law velocity field, rapid acceleration to the terminal velocity (see Figure 2), which can be interpreted as thermal (ram pressure) or cosmic ray (magnetic pressure) driving as each can produce similar rapid-acceleration velocity fields based on the shape of the velocity field (see Figure 1, also 2005ApJ...618..569M; 2016MNRAS.463..541C; Flury2023). Thermal driving occurs when “cool" ($10^{4}\rm~K$) clouds are embedded in an adiabatically expanding hot ($10^{6}\rm~K$) gas (2005ApJ...618..569M). Cosmic ray occurs driving where clouds are accelerated by the diffusion (Breitschwerdt1993; Quataert2022a) or streaming (Ipavich1975; Breitschwerdt1991; Quataert2022b) of relativistic charged particles–launched by hot ionized media such as shocks or supernovae (Colgate1979; Blandford1980)–as they interact with magnetic fields in the ISM, resulting in a radially outward pressure gradient. NGC 2992 is known to contain $\approx 0.5$ kpc radio lobes forming an hourglass or “figure eight” shape, indicative of bi-directional bubbles of hot gas driven by nuclear activity (Wehrle1988; Allen1999). Combining optical and radio measurements indicates the presence of hot cavities which are heated (likely through shocks) by jets or accretion disk winds and within which are entrained clouds of cooler gas (Allen1999; Veilleux2001). The presence of such lobes and the associated hollow cavities suggests either or both (a plausible scenario, e.g., Breitschwerdt2002; Everett2008) mechanisms are involved in launching the outflow in NGC 2992.
@@
Spatial analysis of the outflow in Veilleux2001 found an inclination of $i=68\pm 3^{\circ}$ and an opening angle of $\theta_{o}=62.5-67.5^{\circ}$. OutLines finds $i=53.19\pm 1.32^{\circ}$ and $\theta_{o}=68.60\pm 1.26^{\circ}$. While $\theta_{o}$ is in apparent agreement, ii is significantly discrepant, requiring further assessment of the fit results. To test the reliability of the geometry inferred from the OutLines model, we project the best-fit bidirectional cones onto the plane of the sky and compare with the spatially resolved continuum-subtracted [O iii] doublet flux in the WiFES IFU in Figure 20. We find good agreement between the cone geometry implied by the integrated line profile using OutLines and the observed flux distribution, indicating that modeling with OutLines can recover the spatial geometry without any knowledge of that geometry apriori. Not only have we accurately recovered the spatial flux distribution with OutLines, but we have also inferred the presence of a cavity comprising the inner two thirds of the outflow cone ($\theta_{c}=48.64\pm 1.82^{\circ}$), suggesting that cooler line-emitting clouds can only survive towards the edges of the hot cavity where the thermal wind may be somewhat cooler as it expands and contacts the surrounding ISM.
@@
Figure 20: Greyscale map across the central kiloparsec of NGC 2992 of the continuum-subtracted, integrated [O iii] doublet flux in the S7 DR2 (S7DR2). Drawn cones indicate the geometry inferred from the best-fit OutLines profile to the integrated emission line profile. No spatial information was used to determine the shape of the projected cones. Blue indicates the anterior cone, red indicates the posterior cone. Indigo indicates the disk, here with arbitrary radius scaled for visualization. A PA of 75 was assumed to rotate the projected geometry to the PA of the observation.
@@
With the best-fit parameters, OutLines computes high mass loss rates of at least 40 $\rm M_{\odot}~yr^{-1}$. From the $v_{0}\sim 300$ $\rm km~s^{-1}$ rotational velocity of the galaxy (Veilleux2001), the outflow must be capable of ejecting material from the host galaxy with the characteristic velocity exceeding the escape velocity by a factor of 3. Together, these results suggest NGC 2992 is in a state of AGN-triggered inside-out quenching with the combined effects of hot cavities and the evacuation of cool gas away from the plane of the galaxy, both effects preventing future star formation. Interestingly, this quenching feedback is not sufficiently columnated to be driven by a jet but instead is most likely due to expanding gas heated by winds launched from the accretion disk. While the implications of quenching are clear for galaxy evolution, the association of hot winds and shocks also has implications for the measurement of chemical abundances in AGN (e.g., Dors2021). Here, inferring underlying physics, geometry, and outflow strength from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in the context of AGN feedback.
@@ 5 Conclusion
  • Young (<4<4 Myr) H ii regions undergoing expansion can exhibit two orders of magnitude in gas compression due to wind-triggered shocks even while optically thick radiation pressure is the main driving mechanism of the bulk motion of the gas. Thus, while shocks may dominate the radio continuum, it is radiation pressure from stars which primarily shapes the evolution of H ii regions.
@@ Acknowledgements
We thank the referee for comments which improved this manuscript. We thank the beta testers, including K. Z. Arellano Cordova, K. Duncan, and T. Horn, whose efforts helped to optimize the use of this code. We thank K. Z. Arellano Cordova, C. Esteban, and J. Garcia-Rojas for providing access to Knot A extractions from their reduced MEGARA data. We thank K. Parker for sharing results from the recently accepted Parker2025. We thank K. Duncan for insight regarding organization of this manuscript. We also thank M. Divakara and G. Romero-Cruz for insightful discussions about outflow geometry at the National Institute of Astrophysics, Optics and Electronics (INAOE) and J. Chisholm, C. Martin, and N. Murray for insightful discussions about winds, massive stars, feedback, and HST/COS at the Aspen Center for Physics.
@@
Software, repositories, and other digital resources: corner (corner), emcee (emcee), matplotlib (matplotlib), MAST, NIST/ASD (NIST_ASD), numpy (numpy), OutLines (this work, see also 10.5281/zenodo.11238265), PyNeb (2015A&A...573A..42L), S7 DR2 (S7DR2), scipy (scipy), vygrboi
@@ Appendix A Geometry of Directed Outflows
A.1 Spherical Cap Projection
The ellipse is the projected spherical cap with axes $g$ and $h$ and is offset from the projected outflow source by some distance $S$. These quantities can be expressed in terms of the inclination and opening angle of the ouflow center such that
@@ A.1 Spherical Cap Projection
In the case where $i+\theta>90^{\circ}$, the projection is no longer an ellipse but rather an ellipse plus a lune to account for the projection of the “top” of the spherical cap and an additional lune to the opposing side to account for the projection of the second cone. We visualize such a scenario in the last two panels of Figure 21.
@@ A.2 Cone Inscription of Deprojected Velocities
The fractional arc length $\ell_{\rm w}$ of the deprojected velocity inscribed by the cone can be determined by solving the system of linear equations for a circle and the projected spherical cap area to determine where they intersect, which in turn determine the bounds on the integral(s) over the azimuthal coordinate of the solid angle. The azimuthal “ring” of gas with velocity $w$ is defined by the sine of the angle between the observed and deprojected velocities $u$ and $w$, respectively, such that its radius is
$r_{\rm w}=\sqrt{1-(u/w)^{2}}.$ (44)
Solving the system of equations of the projected cap and the deprojected velocity ring gives intersection points of ($p$, $q$) such that
$p=\pm\sqrt{r_{\rm b}^{2}-q^{2}}$ (45)
@@
$C$ $=r_{\rm b}^{2}+\left(\frac{gS}{h}\right)^{2}-g^{2}.$ (49)
@@
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>0\\ \pi-\arccos{\frac{q}{r_{\rm b}}}&q<0\end{cases}$ (50)
which notably avoids the additional step of calculating $p$. If multiple real solutions for ($p$, $q$) exist due to the lune from the projection of the second cone into the plane, then OutLines simply sums the two arcs.
@@
Figure 21: Directed outflows (left of each panel) projected onto the plane of the sky (right of each panel) for various combinations of $i$ and $\theta_{o}$ (orange). Circles indicate deprojected velocities of u/w=u/w=0.4 (green), 0.6 (blue), and 0.8 (purple). Opaque segments indicate where a deprojected velocity is inscribed by the projected spherical cap. Arrow in the left of each panel indicates direction of the projections onto the plane of the sky (i.e., the line of sight).

### v2

@@ 4.3 Outflow Formation and Blowout in a Starburst Galaxy
It may be tempting to conclude that hot gas or cosmic rays from these supernovae are driving the shells’ expansion, especially since, for the lowest velocity shell at 188 $\rm km~s^{-1}$, the dynamical time to traverse the inner 500 pc is $\la 2.5$ Myr, well after the onset of CCSNe for a burst at least 10 Myr old. That being said, the age reported by Parker et al. (2025) from fits to the UV continuum that Mrk 1486 corresponds to a combination of a substantial young ($\leq 5$ Myr) stellar population and a far older (40 Myr) population. In light of the more detailed star formation history, while CCSNe may be the progenitors of the material in these expanding shells, they may not necessarily be responsible for their acceleration. Here, the shape of the velocity field from OutLines can provide insight into the driving mechanism. Indeed, the velocity index of $\beta=0.7\pm 0.2$ disfavors (but cannot fully not reject) the rapid acceleration associated with ram pressure driving ($\beta\la 0.5$, e.g., Murray et al., 2005; Chisholm et al., 2016). However, these $\alpha$-rich shells are more likely continuously accelerated by optically thin line-driven radiation pressure associated with $\beta=0.8$ (e.g., Pauldrach et al., 1986). Moreover, the total energy injection rate of $10^{40}\rm~erg~s^{-1}$ for the series of shells is roughly consistent with stellar winds, which have $L_{mech}\approx 10^{39}\rm~erg~s^{-1}$ for a given stellar population (see, e.g., Starburst99 or BPASS, Leitherer et al., 2014; Byrne et al., 2022, respectively). Radiative driving by young stellar populations may also account for the less pronounced C ii troughs via ionizing feedback. Furthermore, given that the travel time from the nucleus to a 1 kpc radius (roughly the COS aperture at the 139.1 Mpc distance of Mrk 1486) is $\la 5$ Myr, the >40>40 Myr stellar populations are unlikely to be responsible for these superbubbles. Corroborating evidence for optically thick radiation pressure is the presence of the fluorescent $\lambda 1194,1197$ lines but the absence of resonant line infilling. Such a scenario occurs in the optically thick line limit where photon escape primarily occurs through the fluorescent channel (e.g., Scarlata and Panagia, 2015, their Figures 5-6).
@@ 4.4 Thermal Winds in AGN Radio Lobes
The Seyfert 2 galaxy NGC 2992 has long been known to exhibit conical outflows (Allen et al., 1999; Veilleux et al., 2001). Using archival observations from the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, Thomas et al., 2017) made using WiFES (Dopita et al., 2010), we have confirmed broad, asymmetric [O iii] $\lambda\lambda 4959,5007$ lines after summing over all the spaxels to boost signal to noise in the line profile and to mitigate any aperture effects. The low resolution ($R\approx$ 3 000 or FWHM of 105 $\rm km~s^{-1}$) requires corrections for the LSF by convolving with a Gaussian kernel with $\sigma_{\rm LSF}=$45 $\rm km~s^{-1}$; however, we note that the LSF predominantly impacts the narrow Gaussian core. The high signal in the wings ($S/N\approx 3$ in the blue broad extent of [O iii] $\lambda 4959$) might allow for tight constraints on the wind properties even without LSF corrections – indeed fitting with and without LSF corrections yields similar results for the OutLines parameters. Nevertheless, incorporating the LSF into the forward modeling with OutLines is essential to remain consistent with the observations. As a part of the LSF correction, we fold in the summed best-fit stellar continuum included in S7 to account for stellar contributions to the continuum, including any contamination of the line profile.
@@
The bright intermediate portion of the wings suggests the acceleration power law for the velocity field. Visual assessment of the highly asymmetric line profile in the spectrum of NGC 2992 and comparison with models in Figures 11, 12, and 14 also suggests a hollow cone geometry with an obstructing disk. As such, we subsequently model the [O iii] doublet using OutLines assuming an acceleration power law, an open cone geometry, and an obstructing disk. We impose the following into the prior: line fluxes within 5% of the [O iii] 5007/4959 flux ratio of 2.98 given by atomic data (Storey and Zeippen, 2000) and $\theta_{o}-\theta_{c}>5^{\circ}$. With line profiles predicted by OutLines added to the continuum and subsequently convolved with the WiFES LSF (Dopita et al., 2010), we use emcee to sample the posterior on the outflow parameters. Due to the large number of parameters, potential degeneracies, and sensitivity to initial conditions, we run the sampler four times longer, for a total of 40 000 samples post burn-in. Implementation of our modeling is demonstrated for reference in Appendix C with the best-fit profile shown in Figure 20 and results listed in Table 1. We note that the posterior sampling ultimately rejects the priors on the angles as the best-guess parameters fell within a local minimum.
@@
Figure 20: MCMC fit of an OutLines.Nebular model to the [O iii] $\lambda\lambda 4959,5007$ doublet observed in the nucleus of Seyfert galaxy NGC 2992 as part of the Siding Spring Seyfert Spectroscopic Snapshot Survey (S7, Thomas et al., 2017) DR2, assuming a hollow cone plus disk outflow geometry. Line styles and colors as in Figure 17. Blue line indicates the profile after LSF correction.
@@
The slope $\beta=2.02$ indicates, for the acceleration power law velocity field, rapid acceleration to the terminal velocity (see Figure 2), which can be interpreted as thermal (ram pressure) or cosmic ray (magnetic pressure) driving as each can produce similar rapid-acceleration velocity fields based on the shape of the velocity field (see Figure 1, also Murray et al., 2005; Chisholm et al., 2016; Flury et al., 2023). Thermal driving occurs when “cool" ($10^{4}\rm~K$) clouds are embedded in an adiabatically expanding hot ($10^{6}\rm~K$) gas (Murray et al., 2005). Cosmic ray occurs driving where clouds are accelerated by the diffusion (Breitschwerdt et al., 1993; Quataert et al., 2022b) or streaming (Ipavich, 1975; Breitschwerdt et al., 1991; Quataert et al., 2022a) of relativistic charged particles–launched by hot ionized media such as shocks or supernovae (Colgate and Petschek, 1979; Blandford and Ostriker, 1980)–as they interact with magnetic fields in the ISM, resulting in a radially outward pressure gradient. NGC 2992 is known to contain $\approx 0.5$ kpc radio lobes forming an hourglass or “figure eight” shape, indicative of bi-directional bubbles of hot gas driven by nuclear activity (Wehrle and Morris, 1988; Allen et al., 1999). Combining optical and radio measurements indicates the presence of hot cavities which are heated (likely through shocks) by jets or accretion disk winds and within which are entrained clouds of cooler gas (Allen et al., 1999; Veilleux et al., 2001). The presence of such lobes and the associated hollow cavities suggests either or both (a plausible scenario, e.g., Breitschwerdt et al., 2002; Everett et al., 2008) mechanisms are involved in launching the outflow in NGC 2992.
@@
Spatial analysis of the outflow in Veilleux et al. (2001) found an inclination of $i=68\pm 3^{\circ}$ and an opening angle of $\theta_{o}=62.5-67.5^{\circ}$. OutLines finds $i=53.19\pm 1.32^{\circ}$ and $\theta_{o}=68.60\pm 1.26^{\circ}$. While $\theta_{o}$ is in apparent agreement, ii is significantly discrepant, requiring further assessment of the fit results. To test the reliability of the geometry inferred from the OutLines model, we project the best-fit bidirectional cones onto the plane of the sky and compare with the spatially resolved continuum-subtracted [O iii] doublet flux in the WiFES IFU in Figure 21. We find good agreement between the cone geometry implied by the integrated line profile using OutLines and the observed flux distribution, indicating that modeling with OutLines can recover the spatial geometry without any knowledge of that geometry apriori. Not only have we accurately recovered the spatial flux distribution with OutLines, but we have also inferred the presence of a cavity comprising the inner two thirds of the outflow cone ($\theta_{c}=48.64\pm 1.82^{\circ}$), suggesting that cooler line-emitting clouds can only survive towards the edges of the hot cavity where the thermal wind may be somewhat cooler as it expands and contacts the surrounding ISM.
@@
Figure 21: Greyscale map across the central kiloparsec of NGC 2992 of the continuum-subtracted, integrated [O iii] doublet flux in the S7 DR2 (Thomas et al., 2017). Drawn cones indicate the geometry inferred from the best-fit OutLines profile to the integrated emission line profile. No spatial information was used to determine the shape of the projected cones. Blue indicates the anterior cone, red indicates the posterior cone. Indigo indicates the disk, here with arbitrary radius scaled for visualization. A PA of 75 was assumed to rotate the projected geometry to the PA of the observation.
@@
With the best-fit parameters, OutLines computes high mass loss rates of at least 40 $\rm M_{\odot}~yr^{-1}$. From the $v_{0}\sim 300$ $\rm km~s^{-1}$ rotational velocity of the galaxy (Veilleux et al., 2001), the outflow must be capable of ejecting material from the host galaxy with the characteristic velocity exceeding the escape velocity by a factor of 3. Together, these results suggest NGC 2992 is in a state of AGN-triggered inside-out quenching with the combined effects of hot cavities and the evacuation of cool gas away from the plane of the galaxy, both effects preventing future star formation. Interestingly, this quenching feedback is not sufficiently columnated to be driven by a jet but instead is most likely due to expanding gas heated by winds launched from the accretion disk. While the implications of quenching are clear for galaxy evolution, the association of hot winds and shocks also has implications for the measurement of chemical abundances in AGN (e.g., Dors et al., 2021). Here, inferring underlying physics, geometry, and outflow strength from a single-aperture spectrum demonstrates the power and effectiveness of OutLines in the context of AGN feedback.
@@ 5 Conclusion
  • Young (<4<4 Myr) H ii regions undergoing expansion can exhibit two orders of magnitude in gas compression due to wind-triggered shocks even while optically thick radiation pressure is the main driving mechanism of the bulk motion of the gas. Thus, while shocks may dominate the radio continuum, it is radiation pressure from stars which primarily shapes the evolution of young H ii regions.
@@ Acknowledgements
We thank the beta testers, including K. Z. Arellano Cordova, K. Duncan, A. Hall, and T. Horn, whose efforts helped to optimize the use of this code. We thank K. Z. Arellano Cordova, C. Esteban, and J. Garcia-Rojas for providing access to Knot A extractions from their reduced MEGARA data. We thank K. Parker for sharing results from the recently accepted Parker et al. (2025). We thank K. Duncan for insight and comments regarding this manuscript. We thank A. Rankine and A. Lawrence for suggestions regarding the relativistic corrections. We also thank M. Divakara and G. Romero-Cruz for insightful discussions about outflow geometry at the National Institute of Astrophysics, Optics and Electronics (INAOE) and J. Chisholm, C. Martin, and N. Murray for insightful discussions about winds, massive stars, feedback, and HST/COS at the Aspen Center for Physics.
@@
Software, repositories, and other digital resources: corner (Foreman-Mackey, 2016), curlyBrace, emcee (Foreman-Mackey et al., 2013), matplotlib (Hunter, 2007), MAST, NIST/ASD (Kramida et al., 2024), numpy (van der Walt et al., 2011), OutLines (this work, see also 10.5281/zenodo.11238265), PyNeb (Luridiana et al., 2015), S7 DR2 (Thomas et al., 2017), scipy (P. Virtanen, 2020), vygrboi
@@ Appendix A Geometry of Directed Outflows
A.1 Cones As A Spherical Cap Projection
An outflowing cone on the plane of the sky appears as a the projected spherical cap with axes $g$ and $h$ and is offset from the projected outflow source by some distance $S$. These quantities can be expressed in terms of the inclination and opening angle of the ouflow center such that
@@ A.1 Cones As A Spherical Cap Projection
We illustrate a single case of such a projection in Figure 22. In the case where $i+\theta>90^{\circ}$, the projection is no longer an ellipse but rather an ellipse plus a lune to account for the projection of the “top” of the spherical cap and an additional lune to the opposing side to account for the projection of the second cone. We visualize such a scenario in the last two panels of Figure 23.
Figure 22: A cone geometry in the case of $i=40^{\circ}$, $\theta=30^{\circ}$. Left panels illustrate the cones seen edge-on, i.e., transverse to the line of sight. Right panels illustrate the cones seen face-on, i.e., projected onto the plane of the sky. Solid ellipses indicates the cone projected anterior to source relative to the observe while dashed ellipses indicate the posterior projection. Top panels illustrate the geometric terms $S$, $g$, and $h$ which translate the opening angle $\theta$ and inclination $i$ of the cones into a projection onto the sky. Bottom panels illustrate how an observed velocity $u$ de-projects into an azimuthal band with radius $r_{w}$ with unique azimuthal coverage $\pi\ell$ within the cone defined by the intersection points $p,q$ of the velocity band with the projected cone. High $u/w$ (low $r_{w}$) corresponds to wings of a line profile far from the systemic velocity while low $u/w$ (high $r_{w}$) corresponds to the core of the line profile close to the systemic velocity.
@@ A.2 Cone Inscription of Deprojected Velocities
The fractional arc length $\ell_{\rm w}$ of the deprojected velocity inscribed by the cone can be determined by solving the system of linear equations for a circle and the projected spherical cap area to determine where they intersect, which in turn determine the bounds on the integral(s) over the azimuthal coordinate of the solid angle. The azimuthal “ring” of gas with velocity $w$ has a radius $r_{\rm w}$ defined by the sine of the angle $\varrho$ between the observed and deprojected velocities $u$ and $w$, respectively. Relativistic aberration (i.e., beaming effect) is invariant across the azimuthal angle, ergo a classical treatment in the source frame is still accurate for deriving the geometry of the projected cone. The apparent deprojection radius $r_{\rm w}$ for each azimuthal ring is thus
$r_{\rm w}=\sqrt{1-\left(\frac{u}{w}\right)^{2}},$ (48)
which we show in Figure 22. Solving the system of equations of the projected cap and the deprojected velocity ring gives intersection points of ($p$, $q$) such that
$p=\pm\sqrt{r_{\rm w}^{2}-q^{2}}$ (49)
@@
$C$ $=r_{\rm w}^{2}+\left(\frac{gS}{h}\right)^{2}-g^{2}.$ (53)
@@
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm w}}}&q>0\\ \pi-\arccos{\frac{q}{r_{\rm w}}}&q<0\end{cases}$ (54)
which notably avoids the additional step of calculating $p$. We illustrate such a solution for $\pi\ell$, including the velocity dependence, in Figure 22. If multiple real solutions for ($p$, $q$) exist due to the lune from the projection of the second cone into the plane, then OutLines simply sums the two arcs.
@@
Figure 23: Following Figure 14, we illustrate bidirectional cone geometries (left of each panel) projected onto the plane of the sky (right of each panel) for various combinations of $i$ and $\theta_{o}$ (orange). Circles indicate deprojected velocities of u/w=u/w=0.4 (green), 0.6 (blue), and 0.8 (purple). Opaque segments indicate where a deprojected velocity is inscribed by the projected spherical cap. Arrow in the left of each panel indicates direction of the projections onto the plane of the sky (i.e., the line of sight).